OpenStudy (anonymous):

Subtract. 7/16 – 5/16 A.1/8 B.1/4 C.12/32 D.3/4

OpenStudy (mrnood):

7 (sixteenths) - 5 (sixteenths) = how many sixteenths?

OpenStudy (anonymous):

its a

OpenStudy (anonymous):

2/16

OpenStudy (anonymous):

2/16 divide by 2

OpenStudy (mrnood):

@madib12 it is against the rules to give direct answer but worse than that it DOES NOT help the poster How do they know oyu are right? how will they do the next similar question?

OpenStudy (anonymous):

now simplify

OpenStudy (anonymous):

okay sorry im new I wont do it again

OpenStudy (mrnood):

@prinsess OK you have 2/16 so simplify by dividing top and bottom by 2...

OpenStudy (anonymous):

i got A

OpenStudy (anonymous):

cause 2/2=1 16/2=8

OpenStudy (anonymous):

1/8

OpenStudy (anonymous):

A

OpenStudy (mrnood):

@Trincent it is against the rules to give direct answer but worse than that it DOES NOT help the poster How do they know oyu are right? how will they do the next similar question?

OpenStudy (anonymous):

i was only saying the answer because others had already explained the answer, there for giving the client confidence in the answer. why repeat whats already been said?